OpenStudy (milo123):

Part A: The sun produces 3.9 ⋅ 10^33 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 1.55 ⋅ 10^7 seconds? (5 points) Part B: Which is the more reasonable measurement of the distance between the tracks on a CD: 1.6 ⋅ 10^−3 mm or 1.6 ⋅ 10^3 mm? Justify your answer. (5 points)

OpenStudy (milo123):

@Teddyiswatshecallsme @perl

OpenStudy (mhchen):

In order to help guide and walk you through this problem, think of how to get radiants of energy with the sun's production of radiant energy rate. What is the rate formula? How do you find the total output with the rate formula? By not telling you the direct answer and helping give you hints to walk you through this problem, you will achieve your goal of understanding how this problem works.

OpenStudy (milo123):

( 3.9 x 10^33 ) x ( 1.55 x 10^7 ) = ( 3.9 x 1.55 ) x 10^( 33 + 7 ) = 6.045 x 10^40 So 6.045 x 10^40. ? @mhchen

OpenStudy (wolf1728):

I'd say you did that correctly

OpenStudy (milo123):

sure? lol

OpenStudy (wolf1728):

Yes Also why the measure of time 1.55 *10^7? That's roughly a half year.

OpenStudy (milo123):

ok thanks

OpenStudy (milo123):

can you help me check my answers..theres three

OpenStudy (wolf1728):

Okay

OpenStudy (wolf1728):

Answers to question 1 are correct Going to check question 2

OpenStudy (milo123):

ok

OpenStudy (milo123):

hows the second one?

OpenStudy (wolf1728):

I get an answer of 4 I) (3^8 * 2^-5 *9^0)^-2 = (205)^-2 = 1/(205)^2 = 2.3788^-5 II) (2^-2/3^3)^4 = (.25/27)^4 = (0.0092592593)^4 = 7.3503^-9 III) 3^38 = 2.2877^13 I * II * III = =2.3788^-5 * 7.3503^-9 * 2.2877^13 = 4 Correct again !!

OpenStudy (milo123):

I need help with part b of the last one though....maybe 1.6 x 10^-3?

OpenStudy (wolf1728):

Yes 1.6 * 10^-3 If the spacing of the tracks were 1.6*10^3mm that would mean the space between each track would be 1.6 METERS which is much larger than a CD.

OpenStudy (milo123):

ok

OpenStudy (wolf1728):

Glad I could help

OpenStudy (milo123):

thanks!!:)

OpenStudy (milo123):

wait is 3 right lol i forgot that one:)